OpenStudy (anonymous):

Julianne wants to take a trip around the world. She plans to deposit $175 at the beginning of each month into an investment with a 3.25% interest rate, compounded monthly. How much will she have in the account after 12 years? $30,770.00 $30,853.33 $533,286.11 $550,617.91

OpenStudy (anonymous):

i keep doing the calculations and im not coming up with any of the ANSWERS please help somebody

OpenStudy (anonymous):

the formula for compound interest is : I=P(1+r/4)4n−P where P is the principal amount, n= time period in years and r= rate of interest.

OpenStudy (destinymasha):

it could be ....68.25

OpenStudy (destinymasha):

I x R x P

OpenStudy (campbell_st):

When the annuity is paid at the beginning of the period the formula is \[A = \frac{M[(1 + r)^n -1]}{r} \times (1 + r)\] in your question M = 175, r = 0.0325 and n = 144 substitute and evaluate for the answer.

OpenStudy (anonymous):

$550,617.91 is not the right answer